# SCRREAM : SCan, Register, REnder And Map: A Framework for Annotating Accurate and Dense 3D Indoor Scenes with a Benchmark

### v1

@@ Abstract
Traditionally, 3d indoor datasets have generally prioritized scale over ground-truth accuracy in order to obtain improved generalization. However, using these datasets to evaluate dense geometry tasks, such as depth rendering, can be problematic as the meshes of the dataset are often incomplete and may produce wrong ground truth to evaluate the details. In this paper, we propose SCRREAM, a dataset annotation framework that allows annotation of fully dense meshes of objects in the scene and registers camera poses on the real image sequence, which can produce accurate ground truth for both sparse 3D as well as dense 3D tasks. We show the details of the dataset annotation pipeline and showcase four possible variants of datasets that can be obtained from our framework with example scenes, such as indoor reconstruction and SLAM, scene editing & object removal, human reconstruction and 6d pose estimation. Recent pipelines for indoor reconstruction and SLAM serve as new benchmarks. In contrast to previous indoor dataset, our design allows to evaluate dense geometry tasks on eleven sample scenes against accurately rendered ground truth depth maps.
@@
Due to the flexibility of the approach to scan and annotate scenes for different purposes, we refer to it as a framework. It can be used for various tasks such as indoor reconstruction & SLAM, scene editing & object removal, human reconstruction and 6d pose estimation. In later sections, we detail the steps of the framework and discuss how the scene can be setup to obtain different variants of indoor datasets. To showcase the versatility of the framework we provide eleven example scenes obtained using the proposed framework. The entire dataset is made available for the community [^1] Dataset link and toolkit can be found in https://github.com/Junggy/SCRREAM and we provide a benchmark for 3D geometric tasks using existing state of the art (SoTA) methods for indoor reconstruction & SLAM on our example scenes.
@@ 2 Related Works
Early datasets obtain their annotations by using directly the depth sensor readings [1, 2], where invalid or empty pixels are manually annotated to provide dense 3D data for evaluation [1]. Later, the paradigm for annotation changed, i.e. the indoor scene is reconstructed by fusing a sequence of depth images obtained by multi modal camera trajectories [3, 4, 5] to produce more complete scene geometry. Depth and other 3d information is produced on the image plane by rendering the mesh using the camera pose. To improve the quality and density of the datasets, annotation methods that pre-scan the scene, record [6] and localize the camera are developed [7, 8]. Although this annotation technique produces a significantly better mesh quality of the scene, it suffers from line of sight problems of the scanner as the scanning is done on a scene level. Ensuring every single object in the scene is complete is a non-trivial task. As a result, dense ground truth generated in this way suffers from holes and missing parts. Therefore, scene-level reconstruction methods like neural radiance fields (NeRF) [12] variants or 3D Gaussian Splatting [13] evaluate their geometry only via qualitative evaluation on their depth prediction.
@@ 4 Dataset Variations
Indoor Reconstruction and SLAM Dataset. In this part, we consider indoor reconstruction and SLAM as the major task of our dataset. The dataset can be obtained by following the annotation pipeline described above. Dense depth maps and instance masks are rendered with camera poses and annotated meshes of the scene as shown in Fig. 4 (c). We acquire eleven scenes for this task and provide the benchmarks that evaluate the quality of view synthesis and 3D geometry tasks using recent approaches in Sec. 5.

### v2

@@ Abstract
Traditionally, 3d indoor datasets have generally prioritized scale over ground-truth accuracy in order to obtain improved generalization. However, using these datasets to evaluate dense geometry tasks, such as depth rendering, can be problematic as the meshes of the dataset are often incomplete and may produce wrong ground truth to evaluate the details. In this paper, we propose SCRREAM, a dataset annotation framework that allows annotation of fully dense meshes of objects in the scene and registers camera poses on the real image sequence, which can produce accurate ground truth for both sparse 3D as well as dense 3D tasks. We show the details of the dataset annotation pipeline and showcase four possible variants of datasets that can be obtained from our framework with example scenes, such as indoor reconstruction and SLAM, scene editing & object removal, human reconstruction and 6d pose estimation. Recent pipelines for indoor reconstruction and SLAM serve as new benchmarks. In contrast to previous indoor dataset, our design allows to evaluate dense geometry tasks on eleven sample scenes against accurately rendered ground truth depth maps. (https://sites.google.com/view/scrream/about)
@@
With SCRREAM we follow the spirit of our previous 3D dataset projects PhoCal [10], Hammer [9] and HouseCat6D [11] that democratized advances in this field for smaller scale 3D perception and manipulation tasks by describing all pipeline parts and the hardware components in detail, enabling the community to apply the acquisition principle on their own hardware setups. We further provide the source code of visualization tools with detailed documentation to allow users to utilize the SCRREAM data for downstream tasks. The entire dataset is made available for the community [^1] Dataset link and toolkit can be found in https://github.com/Junggy/SCRREAM and we provide a benchmark for 3D geometric tasks using existing state of the art (SoTA) methods for indoor reconstruction & SLAM on our example scenes.
@@ 2 Related Works
Figure 2: Dataset Geometry Comparison. Our dataset features high-quality, complete meshes of the scene. In comparison, commonly used datasets such as Replica [6] and ScanNet++ [7] suffer from over-smoothed or incomplete meshes (Zoom in for details). We include more in-depth comparisons in the supplementary material pdf and video file.
Early datasets obtain their annotations by using directly the depth sensor readings [1, 2], where invalid or empty pixels are manually annotated to provide dense 3D data for evaluation [1]. Later, the paradigm for annotation changed, i.e. the indoor scene is reconstructed by fusing a sequence of depth images obtained by multi modal camera trajectories [3, 4, 5] to produce more complete scene geometry. Depth and other 3d information is produced on the image plane by rendering the mesh using the camera pose. To improve the quality and density of the datasets, annotation methods that pre-scan the scene, record [6] and localize the camera are developed [7, 8]. Although this annotation technique produces a significantly better mesh quality of the scene, it suffers from line of sight problems of the scanner as the scanning is done on a scene level. Ensuring every single object in the scene is complete is a non-trivial task. As a result, dense ground truth generated in this way suffers from holes and missing parts (Fig. 2). Therefore, scene-level reconstruction methods like neural radiance fields (NeRF) [14] variants or 3D Gaussian Splatting [15] evaluate their geometry only via qualitative evaluation on their depth prediction.
@@ 4 Dataset Variations
Indoor Reconstruction and SLAM Dataset. In this part, we consider indoor reconstruction and SLAM as the major task of our dataset. The dataset can be obtained by following the annotation pipeline described above. Dense depth maps and instance masks are rendered with camera poses and annotated meshes of the scene as shown in Fig. 5 (c). We acquire eleven scenes for this task (see Fig. 10 for more details) and provide the benchmarks that evaluate the quality of view synthesis and 3D geometry tasks using recent approaches in Sec. 5.
@@ 6 Limitation and Conclusion
This work is from R&D Project at TUM funded by Huawei Research & Development (UK) Ltd.
Figure 10: Example of Indoor Reconstruction and SLAM Scenes.